40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 11509     Entry time: Fri Aug 14 23:49:34 2015
Author: Koji 
Type: Summary 
Category: General 
Subject: B&K Shaker fixed 

I fixed a shaker that was claimed to be broken. I had to cut the rubber membrane to open the head.

Once it was opened, the cause of the trouble was obvious. The soldering joint could not put up with the motion of the head.

It is interesting to see that the spring has the damping layer between the metal sheets.

After the repair the DC resistance was measured. It was 1.9Ohm. The side of the shaker chassis said "3.5Ohm, Max 15VA". So it can take more than 4A (wow).

I gave 2A DC from the bench top supply and turn the current on and off. I could confirm the head was moving.

I'll claim the use of this shaker for the seismometer development.

Attachment 1: IMG_1947.JPG  6.076 MB  | Show | Show all
ELOG V3.1.3-